# Stochastic heat flow by moments

### v1

@@ 4.3. Tightness
Put $c=c(n,\varphi,I,p,g,g^{\prime})$, put $\underline{s}:=s\wedge s^{\prime}$, $\overline{s}:=s\vee s^{\prime}$, do similarly for $\underline{t}$ and $\overline{t}$, and write $\|\cdot\|_{2n,\mathbf{P}}:=\mathbf{E}[(\cdot)^{2n}]^{1/2n}$. Adding and subtracting terms and using the triangle inequality give
[…]
$\leq\sum_{u=\underline{t},\overline{t}}\big{\|}\big{(}Z^{\varepsilon}_{\underline{s},u}-Z^{\varepsilon}_{\overline{s},u}\big{)}g\otimes g^{\prime}\big{\|}_{2n,\mathbf{P}}+\sum_{u=\underline{s},\overline{s}}\big{\|}\big{(}Z^{\varepsilon}_{u,\overline{t}}-Z^{\varepsilon}_{u,\underline{t}}\big{)}g\otimes g^{\prime}\big{\|}_{2n,\mathbf{P}}\ .$ (4.33)
Consider the first sum in (4.33). Use (4.15) to write $Z^{\varepsilon}_{\underline{s},u}=Z^{\varepsilon}_{\underline{s},\overline{s}}Z^{\varepsilon}_{\overline{s},u}$, decompose $Z^{\varepsilon}_{\underline{s},\overline{s}}$ into $\mathsf{p}^{\scriptscriptstyle\star}_{\underline{s},\overline{s}}+W^{\varepsilon}_{\underline{s},\overline{s}}$, bound the summand by $\|(Z^{\varepsilon}_{\overline{s},u}-\mathsf{p}^{\scriptscriptstyle\star}_{\underline{s},\overline{s}}Z^{\varepsilon}_{\overline{s},u})g\otimes g^{\prime}\|_{2n,\mathbf{P}}+\|W^{\varepsilon}_{\underline{s},\overline{s}}Z^{\varepsilon}_{\overline{s},u}g\otimes g^{\prime}\|_{2n,\mathbf{P}}$, and use (4.10) and $\overline{s}-\underline{s}=|s-s^{\prime}|$ to evaluate these expectations. Doing so gives
$\big{\|}\big{(}Z^{\varepsilon}_{\underline{s},u}-Z^{\varepsilon}_{\overline{s},u}\big{)}g\otimes g^{\prime}\big{\|}_{2n,\mathbf{P}}\leq$ $\big{\langle}g^{\otimes 2n},\big{(}\mathbf{1}-e^{\frac{1}{2}|s-s^{\prime}|\Delta}\big{)}^{\otimes 2n}\,\mathcal{Q}^{[\![2n]\!],\varepsilon}(u-\overline{s})\,g^{\prime}{}^{\otimes 2n}\big{\rangle}^{1/2n}$ (4.34)
@@
where $\Delta$ denotes the Laplacian on $\mathbb{R}^{2}$. By the Cauchy–Schwarz inequality, the right-hand side of (4.34) is bounded by $\|g-\exp(\frac{1}{2}|s-s^{\prime}|\Delta)g\|_{2}\cdot\|\mathcal{Q}^{[\![2n]\!],\varepsilon}(u-\overline{s})g^{\prime}{}^{\otimes 2n}\|_{2}^{1/2n}$. The second factor is bounded by $c$ thanks to (4.9), while the first factor is bounded by $c\,|s-s^{\prime}|$ because $g\in C_{\mathrm{c}}^{2}(\mathbb{R}^{2})$ (recall that $D\subset C_{\mathrm{c}}^{2}(\mathbb{R}^{2})$). Further, since $I$ is bounded, $|s-s^{\prime}|\leq c\,|s-s^{\prime}|^{\frac{1}{2}(1-\frac{2}{p})}$. The term in (4.35) is bounded by $c\,|s-s^{\prime}|^{\frac{1}{2}(1-\frac{2}{p})}$ thanks to Lemma 4.1 for $a=1$ and (4.9). Applying the same argument to the second sum in (4.33) completes the proof. ∎
@@
The bound (2.10a) follows from Young’s convolution inequality. The bound (2.10b) follows by writing $|f|^{2}=(|f|e^{a|x|_{1}})^{2}\cdot e^{-2a|x|_{1}}$ and applying Hölder’s inequality with exponents $p/2$ and $p/(p-2)$. The bound (2.10c) follows by applying Hölder’s inequality with exponents $q$ and $p$. Next, the equalities in (2.10d)–(2.10e) follows by definition; see (2.5a) and (2.7). To prove the inequalities, put $\alpha=ij$ and write $\langle g,\mathcal{P}^{\omega}_{\alpha}(t)f\rangle$ as
@@
We begin by using the comparison argument in [ST24] to reduce the problem to the $a=0$ case. Consider the multiplicative operator $\mathcal{E}:L^{p}(\mathbb{R}^{2})\to L^{p,a}(\mathbb{R}^{2})$, $\psi\mapsto e^{a|\cdot|_{1}}\psi$ and express the weights in the norm as conjugations:
@@
To tackle the conjugations in (A.3), we use the following bound, which is not hard to check and is proven in [ST24, Equation (2.15)]:
@@
[…]
@@
A discrete version of the right-hand side of (A.7) was considered in [CSZ23a, Lemma 6.8] with $h=|\omega|$ and $|I|=|J|=h-1$. The proof applies also to the right-hand side of (A.7) and bound it by $c\,\|f\|_{q}\,\|g\|_{p}$. This proves (2.9d). Finally, applying [GQT21, Lemma 8.4] and (2.10a) to (A.5e) gives (2.9d). ∎
@@
Insert (A.10)–(A.11) and (2.15) into (A.9), bound the size of $\sum_{\kappa\in I^{\mathrm{c}}}$ by $12m+$ 1, and insert the result into (A.8). We hence bound the left-hand side of (A.8) by
@@
Bounding the last sum by $c_{1}^{m}+mc_{0}^{2}\,(c_{1}+c_{0}^{2}/\log(\lambda-c_{0}-1))^{m-1}$ completes the proof. ∎
@@
Recall that $D=\{g_{D,1},g_{D,2},\ldots\}$, take any $N<\infty$, and consider the projection $\Pi_{ij}:C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},M_{+}(\mathbb{R}^{4}))\to C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$: $\mu\mapsto\mu g_{D,i}\otimes g_{D,j}$. For every $i,j\in\mathbb{Z}_{>0}$, since $\{\Pi_{ij}Z^{\varepsilon}\}_{\varepsilon\in(0,1]}$ is tight, there exists a compact $K_{ij}\subset C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$ such that $\mathbf{P}[\Pi_{ij}Z^{\varepsilon}\notin K_{ij}]<2^{-i-j-N}$, for all $\varepsilon\in(0,1]$. Set $K:=\cap_{i,j=1}^{\infty}\Pi_{ij}^{-1}(K_{ij})$. This set is compact. To see why, take any sequence in $K$, enumerate $\Pi_{ij}$ for $i,j\in\mathbb{Z}_{>0}$, and use the diagonal argument to extract a subsequence $\{\mu_{k}\}_{k}$ such that $\Pi_{ij}\mu_{k}$ converges as $k\to\infty$ for every $i,j$. Referring to (1.6) and (2.22), we see that $\{\mu_{k}\}_{k}$ is Cauchy with respect to $\rho$, so it converges. This proves that $K$ is compact. By the union bound, $\mathbf{P}[Z^{\varepsilon}\notin K]\leq 2^{-N}\sum_{i,j}2^{-i-j}=2^{-N}$, for all $\varepsilon\in(0,1]$, so $\{Z^{\varepsilon}\}_{\varepsilon\in(0,1]}$ is tight. ∎

### v3

@@ 4.3. Tightness
Put $c=c(n,\phi,I,p,g,g^{\prime})$, put $\underline{s}:=s\wedge s^{\prime}$, $\overline{s}:=s\vee s^{\prime}$, do similarly for $\underline{t}$ and $\overline{t}$, and write $\|\cdot\|_{2n,\mathbf{P}}:=\mathbf{E}[(\cdot)^{2n}]^{1/2n}$. Adding and subtracting terms and using the triangle inequality give
$\big\|$ $\big(Z^{\theta,\varepsilon}_{s,t}-Z^{\theta,\varepsilon}_{s^{\prime},t^{\prime}}\big)g\otimes g^{\prime}\big\|_{2n,\mathbf{P}}$ (4.33)
$\leq\sum_{u=\underline{t},\overline{t}}\big\|\big(Z^{\theta,\varepsilon}_{\underline{s},u}-Z^{\theta,\varepsilon}_{\overline{s},u}\big)g\otimes g^{\prime}\big\|_{2n,\mathbf{P}}+\sum_{u=\underline{s},\overline{s}}\big\|\big(Z^{\theta,\varepsilon}_{u,\overline{t}}-Z^{\theta,\varepsilon}_{u,\underline{t}}\big)g\otimes g^{\prime}\big\|_{2n,\mathbf{P}}\ .$ (4.34)
Consider the first sum in (4.34). Use (4.15) to write $Z^{\theta,\varepsilon}_{\underline{s},u}=Z^{\theta,\varepsilon}_{\underline{s},\overline{s}}Z^{\theta,\varepsilon}_{\overline{s},u}$, decompose $Z^{\theta,\varepsilon}_{\underline{s},\overline{s}}$ into $\mathsf{p}^{\scriptscriptstyle\star}_{\underline{s},\overline{s}}+W^{\varepsilon}_{\underline{s},\overline{s}}$, bound the summand by $\|(Z^{\theta,\varepsilon}_{\overline{s},u}-\mathsf{p}^{\scriptscriptstyle\star}_{\underline{s},\overline{s}}Z^{\theta,\varepsilon}_{\overline{s},u})g\otimes g^{\prime}\|_{2n,\mathbf{P}}+\|W^{\varepsilon}_{\underline{s},\overline{s}}Z^{\theta,\varepsilon}_{\overline{s},u}g\otimes g^{\prime}\|_{2n,\mathbf{P}}$, and use (4.10) and $\overline{s}-\underline{s}=|s-s^{\prime}|$ to evaluate these expectations. Doing so gives
$\big\|\big(Z^{\theta,\varepsilon}_{\underline{s},u}-Z^{\theta,\varepsilon}_{\overline{s},u}\big)g\otimes g^{\prime}\big\|_{2n,\mathbf{P}}\leq$ $\big\langle g^{\otimes 2n},\big(\mathbf{1}-e^{\frac{1}{2}|s-s^{\prime}|\Delta}\big)^{\otimes 2n}\,\mathcal{Q}^{[\![2n]\!],\varepsilon}(u-\overline{s})\,g^{\prime}{}^{\otimes 2n}\big\rangle^{1/2n}$ (4.35)
@@
where $\Delta$ denotes the Laplacian on $\mathbb{R}^{2}$. By the Cauchy–Schwarz inequality, the right-hand side of (4.35) is bounded by $\|g-\exp(\frac{1}{2}|s-s^{\prime}|\Delta)g\|_{2}\cdot\|\mathcal{Q}^{[\![2n]\!],\varepsilon}(u-\overline{s})g^{\prime}{}^{\otimes 2n}\|_{2}^{1/2n}$. The second factor is bounded by $c$ thanks to (4.9), while the first factor is bounded by $c\,|s-s^{\prime}|$ because $g\in C_{\mathrm{c}}^{2}(\mathbb{R}^{2})$ (recall that $D\subset C_{\mathrm{c}}^{2}(\mathbb{R}^{2})$). Further, since $I$ is bounded, $|s-s^{\prime}|\leq c\,|s-s^{\prime}|^{\frac{1}{4}(1-\frac{2}{p})}$. The term in (4.36) is bounded by $c\,|s-s^{\prime}|^{\frac{1}{4}(1-\frac{2}{p})}$ thanks to Lemma 4.1 for $a=1$ and (4.9). Applying the same argument to the second sum in (4.34) completes the proof. ∎
@@
The bound (2.10a) follows from Young’s convolution inequality. The bound (2.10b) follows by writing $|f|^{2}=(|f|e^{a|x|_{1}})^{2}\cdot e^{-2a|x|_{1}}$ and applying Hölder’s inequality with exponents $p/2$ and p/(p2)p/(p-2). The bound (2.10c) follows by applying Hölder’s inequality with exponents $q$ and $p$. Next, the equalities in (2.10d)–(2.10e) follows by definition; see (2.5a) and (2.7). To prove the inequalities, put $\alpha=ij$ and write $\langle g,\mathcal{P}^{\omega}_{\alpha}(t)f\rangle$ as
@@
We begin by using the comparison argument in [ST25] to reduce the problem to the $a=0$ case. Consider the multiplicative operator $\mathcal{E}:L^{p}(\mathbb{R}^{2})\to L^{p,a}(\mathbb{R}^{2})$, $\psi\mapsto e^{a|\cdot|_{1}}\psi$ and express the weights in the norm as conjugations:
@@
To tackle the conjugations in (A.3), we use the following bound, which is not hard to check and is proven in [ST25, Equation (2.15)]:
@@
We now complete the proof based on (A.5). First, applying (2.10d) to (A.5a)–(A.5b) gives (2.9a). Turning to (A.5c), we write $\mathcal{P}^{\omega}_{\alpha\alpha^{\prime}}(t)=\mathcal{P}^{\alpha\cup\alpha^{\prime}}_{\alpha\alpha^{\prime}}(t)\otimes\mathcal{P}^{\omega\setminus\alpha}(t)$ and use (2.10a) the bound $\|\mathcal{P}^{\omega\setminus\alpha}(t)\|_{p}$ by $1$. Doing so shows that the right-hand side of (A.5c) is bounded by $c\,e^{ct}\|\mathcal{P}^{\alpha\cup\alpha^{\prime}}_{\alpha\alpha^{\prime}}(t)\|_{p}$. Without loss of generality, consider the cases $(\alpha,\alpha^{\prime})=(12,23)$ and $(\alpha,\alpha^{\prime})=(12,34)$. For the first case, write $\langle|f|,\mathcal{P}^{123}_{12\ 23}(t)|g|\rangle$ explicitly as
@@
A discrete version of the right-hand side of (A.7) was considered in [CSZ23a, Lemma 6.8] with $h=|\omega|$ and $|I|=|J|=|\omega|-1$. The proof applies also to the right-hand side of (A.7) and bound it by $c\,\|f\|_{q}\,\|g\|_{p}$. This proves (2.9d). Finally, applying [GQT21, Lemma 8.4] and (2.10a) to (A.5e) gives (2.9d). ∎
@@
Insert (A.10)–(A.11) and (2.15) into (A.9), bound the size of $\sum_{\kappa\in I^{\mathrm{c}}}$ by $2m+1$, and insert the result into (A.8). We hence bound the left-hand side of (A.8) by
@@
Bounding the last sum by $c_{0}^{m+1}\,(c_{1}^{m}+mc_{0}^{2}\,(c_{1}+c_{0}/\log(\lambda-c_{0}-1))^{m-1})$ completes the proof. ∎
@@
Recall that $D=\{g_{D,1},g_{D,2},\ldots\}$, take any $N<\infty$, and consider the projection $\Pi_{ij}:C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathcal{M}_{+}(\mathbb{R}^{4}))\to C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$: $\mu\mapsto\mu g_{D,i}\otimes g_{D,j}$. For every $i,j\in\mathbb{Z}_{>0}$, since $\{\Pi_{ij}Z^{\theta,\varepsilon}\}_{\varepsilon\in(0,1]}$ is tight, there exists a compact $K_{ij}\subset C(\mathbb{R}^{2}_{\scriptscriptstyle\leq},\mathbb{R})$ such that $\mathbf{P}[\Pi_{ij}Z^{\theta,\varepsilon}\notin K_{ij}]<2^{-i-j-N}$, for all $\varepsilon\in(0,1]$. Set $K:=\cap_{i,j=1}^{\infty}\Pi_{ij}^{-1}(K_{ij})$. This set is compact. To see why, take any sequence in $K$, enumerate $\Pi_{ij}$ for $i,j\in\mathbb{Z}_{>0}$, and use the diagonal argument to extract a subsequence $\{\mu_{k}\}_{k}$ such that $\Pi_{ij}\mu_{k}$ converges as $k\to\infty$ for every $i,j$. Referring to (1.9) and (2.22), we see that $\{\mu_{k}\}_{k}$ is Cauchy with respect to $\rho$, so it converges. This proves that $K$ is compact. By the union bound, $\mathbf{P}[Z^{\theta,\varepsilon}\notin K]\leq 2^{-N}\sum_{i,j}2^{-i-j}=2^{-N}$, for all $\varepsilon\in(0,1]$, so $\{Z^{\theta,\varepsilon}\}_{\varepsilon\in(0,1]}$ is tight. ∎